In dire need of help, I'm drowning and trying to stay afloat.
When using the quadratic formula to solve a quadratic equation (ax2 + bx + c = 0), the discriminant is b2 - 4ac. This discriminant can be positive, zero, or negative.
Create three unique equations where the discriminant is positive, zero, or negative. For each case, explain what this value means to the graph of y = ax2 + bx + c.

Answer by venugopalramana(3286)

THIS HAS BEE ALREADY ANSWERED EARLIER..SEE BELOW
---------------------------------------------------------------
Linear-equations/31502: When using the quadratic formula to solve a quadratic equation (ax2 + bx + c = 0), the discriminant is b2 - 4ac. This discriminant can be positive, zero, or negative.
Create three unique equations where the discriminant is positive, zero, or negative. For each case, explain what this value means to the graph of y = ax2 + bx + c.

1 solutions
Answer 18248 by venugopalramana(1268)on 2006-03-28 05:24:24 
When using the quadratic formula to solve a quadratic equation (ax2 + bx + c = 0), the discriminant is b2 - 4ac. This discriminant can be positive, zero, or negative.
Create three unique equations where the discriminant is positive
Y=X^2-5X+6=0
D=5^2-4*6=25-24=1
Y HAS POSITIVE AND NEGATIVE VALUES
, zero,
Y=X^2-4X+4=0
D=4^2-4*4=0
Y IS ALWAYS >=0
or negative
Y=X^2+4=0
D=0-4*4=-16
Y IS ALWAYS POSITIVE.(OR IT CAN BE ALWAYS NEGATIVE...AS IN CASE OF Y=-X^2)
. For each case, explain what this value means to the graph of y = ax2 + bx + c.
graph( 600, 600, -5, 5, -10, 10, x^2-5*x+6,x^2-4*x+4,x^2+4)